# Quantum circuit complexity and unsupervised machine learning of topological order

### v1

@@ III Unsupervised machine learning of topological order derived from quantum circuit complexity
For a family of smoothly parametrized quantum states $\rho(s)$ by $s\in[0,\,1]$, the quantum Fisher information (QFI) of the quantum state with respect to the parameter $s$ is given by Helstrom (1976); Holevo (1982); Liu et al. (2020b) $\mathcal{F}_{Q}\left(s\right)=\mathrm{tr}\left[\rho\left(s\right)\mathcal{L}^{2}\right]$, with the symmetric logarithmic derivative (SLD) $\mathcal{L}$ determined by $\partial_{s}\rho\left(s\right)=\frac{1}{2}\left[\rho\left(s\right)\mathcal{L}+\mathcal{L}\rho\left(s\right)\right]$. The QFI, which plays a central role in quantum parameter estimation and quantum metrology, can be related to the infinitestmal Bures distance Braunstein and Caves (1994) with
@@
$\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\underset{G(s)}{\mathrm{inf}}\ \frac{1}{2}\int^{1}_{0}\sqrt{{\mathcal{F}}_{Q}(s)}\ \ \mathrm{d}s,$ (7)
which represents the optimal Bures distance between the two quantum states Taddei et al. (2013), reflecting the overall fidelity variation, as indicated by (6). Note that we have a symmetric QFC under permmutations of its inputs, i.e., $\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\mathcal{C}_{\mathcal{F}}\left(\rho_{1}\rightarrow\rho_{0}\right)$.
@@
(Quantum circuit complexity upper bounds fidelity change.)  The QCC for the QPP from $\rho_{0}$ to $\rho_{1}$ is lower bounded by the QFC and the Bures distance,
@@
The proof of Theorem 1 is given in Appendix A. Note that, without the constraint of geometric locality of the quantum circuit, the overall Bures distance in (8) is equivalent to the $L_{2}$-norm distance between the two density matrices, which has been shown not to be a suitable distance measure for topological phases Che et al. (2020).
@@
[Quantum circuit complexity upper bounds entanglement change (A generalized version of Observation 1 in Eisert (2021)).] The QCC for the QPP from $\rho_{0}$ to $\rho_{1}$ with a geometrically local quantum circuit (with respect to the $n$-qubit chain) is lower bounded by
@@
The proof of Theorem 2 is given in Appendix A. Theorem 2 can be readily generalized to situations where the qubits are arranged in higher spatial dimensions and other geometries, provided that the quantum circuit is applied in a geometrically local manner accordingly. Nielsen’s QCC is then lower bounded by the entanglement change distributed over suitable cuts covering non-trivial bonds of the system.
@@
IV Relation to classical shadows
@@ V Numerical experiments
V.1 The bond-alternating XXZ spin chain
@@
In Fig. 2, we present a comparison between the fidelity- and entanglement-based kernels for this model, respectively, together with the nonlinear embedding into a two-dimensional representation space from the diffusion map algorithm Coifman et al. (2005); Nadler et al. (2006); Rodriguez-Nieva and Scheurer (2019); Che et al. (2020); Scheurer and Slager (2020). The dataset consists of ground-state properties (fidelities and entanglements) obtained from DMRG calculations of the model. Fixing $h_{0}=0$, $\delta=3$ and varying the value of $J_{2}/J_{1}$, we find that the unsupervised learning produces distinct clusters that can be associated with three phases (trivial, symmetry broken, and topological), in good agreement with independent calculations of topological invariants Pollmann and Turner (2012); Elben et al. (2020); Huang et al. (2022). More details can be found in the caption of Fig. 2. While we are ultimately aiming at the unsupervised learning of topological order in the presence of non-local quantum entanglement, Fig. 2 shows that the fidelity- and entanglement-based kernels also perform well when clustering symmetry-broken phases and symmetry-protected topological orders Elben et al. (2020) that are characterized by short-range entanglement.
@@
V.2 Kitaev’s toric code
@@
where $v$ denotes a vertex and $p$ is the face of the lattice, respectively; $A_{v}=\underset{j\in star(v)}{\prod}X_{j}$ and $B_{p}=\underset{j\in boundary(p)}{\prod}Z_{j}$ are the Kitaev’s star and face operators, respectively. The number of qubits for this model is $n=2L_{x}L_{y}$. The toric code ground state is gapped and exhibits area-law entanglement Kitaev and Preskill (2006). Therefore, for the purpose of a proof-of-principle demonstration and in order to obtain the entanglement profile for the kernel with a small system size, we use a MPS representation and DMRG to find a reasonably accurate solution for the ground state. For instance, numerical experiments show that, for a $4\times 4$ lattice with periodic boundary conditions in both dimensions, a maximal MPS bond dimension of $64$ is sufficient to produce accurate results.
@@
In Fig. 3 we plot a comparison between the unsupervised machine learning of the toric code ground state (blue dots) and the random product states (RPS) (red dots) of qubits (i.e., random bit strings of spin ups and downs), with fidelity- and entanglement-based kernels, respectively. Both kernels perform well in clustering the topologically quantum ordered state and the trivial phase, while the entanglement-based one is more robust under application of two-qubit random unitaries (with respect to the Haar measure) to both the toric code and the RPS. Figure 3 (a) is similar to the result obtained from the shadow kernel learning in Huang et al. (2022), confirming the relation to classical shadows that was elaborated above. The entanglement created by applying the two-qubit random circuit grows exponentially with the circuit depth. More details about the plot can be found in the caption of Fig. 3.
V.3 The extended toric code
@@
VI Discussion and outlook
In summary, we used quantum circuit complexity, a vital concept in quantum computation, to understand and to build viable distance measures and kernels for the unsupervised machine learning of topological order, targeting interpretability and generalizability. By resorting to Nielsen’s quantum circuit complexity, two easier-to-implement kernels based on fidelity and entanglement were proposed and their effectiveness numerically verified.
The shadow kernel learning of Huang et al. (2022) results as a natural derivative of our method based on the Nielsen’s QCC, when the fidelity-based similarity and classical shadow representation of the quantum state are used. The alternative kernel based on the entanglement entropy profile encodes more information about entanglement patterns in topological orders, and thus exhibits superior robustness against two-qubit random noises; moreover, they are more interpretable in terms of the relationship between topological quantum orders and long-range entanglement. Recent theoretical studies Vermersch et al. (2024a, b) have proposed protocols for the efficient measurement of many-body entropies, which indicates that the entanglement-based kernel can be calculated efficiently from experimental data. This implies that our method can be readily applied to experimentally prepared multi-qubit quantum states; for instance, data obtained from near-term quantum computing hardware. When limited to SPT order and compared to the recent results in Wu et al. (2024) about neural-network learning topological orders of the bound-alternating XXZ model, our method is fully unsupervised and results in sharper phase boundaries in clustering quantum phases of this model while providing good interpretability.
@@ VI Discussion and outlook
VII Acknowledgments
@@
Appendix A Proof of Theorems
@@
${\cal{F}}_{Q}(s)$ $=$ $4\left(\langle\partial_{s}\psi(s)|\partial_{s}\psi(s)\rangle-\left|\langle\partial_{s}\psi(s)|\psi(s)\rangle\right|^{2}\right)$ (20)
@@
$\text{Var}\left[G(s)\right]\leq\frac{1}{4}\left[\lambda_{\text{max}}(s)-\lambda_{\text{min}}(s)\right]^{2}\leq\|G(s)\|^{2}_{\infty},$ (21)
where $\lambda(s)$ is the eigenvalue of $G(s)$; and to obtain the last inequality, we have used $|a-b|\leq|a|+|b|$ for real numbers $a$ and $b$, and that, by definition, the operator norm of the generator $\|G(s)\|_{\infty}=|\lambda(s)|_{\text{max}}$. The first inequality in (21) is saturated when $|\psi(s)\rangle=\frac{1}{\sqrt{2}}\left(|\lambda_{\text{max}}(s)\rangle+|\lambda_{\text{min}}(s)\rangle\right)$, i.e., the quantum state is an equal weight superposition of eigenstates with maximal and minimum eigenvalues of $G(s)$, respectively; and the last inequality is saturated when $\lambda_{\text{max}}(s)=-\lambda_{\text{min}}(s)>0$.
@@
$\|G(s)\|_{\infty}\leq\sum_{\sigma}|h_{\sigma}(s)|,$ (22)
@@
$\sqrt{\mathcal{F}_{Q}(s)}\leq 2\sum_{\sigma}|h_{\sigma}(s)|$ (23)
for all possible paths $h_{\sigma}(s)$ ($s\in[0,1]$). Then by integrating over $s\in[0,1]$ on both sides of (23) followed by taking the infimum over all possible paths, the first inequality in (8) holds naturally, i.e.,
$\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)\leq{\cal{C}}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right).$ (24)
The second inequality in (8),
$\frac{D_{B}\left(\rho_{0},\rho_{1}\right)}{\sqrt{2}}$ $\leq$ ${\cal{C}}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right),$ (25)
@@
$G(s)=\sum_{i}h_{i}(s)\ O_{i},$ (26)
where $O_{i}$ is an orthonormal operator basis with operator norm $\|O_{i}\|_{\infty}\leq 1$, then, by replacing the summation in the right-hand side of (23) to go over all orthonormal operator bases of the generator, (23) also holds, i.e.,
$\sqrt{\mathcal{F}_{Q}(s)}\leq 2\sum_{i}|h_{i}(s)|.$ (27)
Next, a suitable similarity measure for topological quantum order imposes the additional constraint of geometric locality of the QPP, which requires that $O_{i}$ is a geometrically local operator for all possible values of $i$, i.e., $O_{i}$ is supported on a constant number of neighboring particles. We consider a reduced density matrix $\rho\left(s|\Delta\right)$ supported on a subsystem $\Delta$ of constant size, undergoing a nonunitary evolution along the unitary QPP of the total system generated by $G(s)$. While the relation between the Bures distance and the QFI in (6) still holds for mixed quantum states with the Uhlmann-Jozsa fidelity, the expression of the QFI for $\rho\left(s|\Delta\right)$ with respect to $s$ is more complicated than (20). Now the QFI has the general form given by the symmetric logarithmic derivative (SLD) $\mathcal{L}$, with
$\mathcal{F}_{Q}\left(s|\Delta\right)=\mathrm{tr}\left[\rho\left(s|\Delta\right)\mathcal{L}^{2}\right],$ (28)
@@
$\partial_{s}\rho\left(s|\Delta\right)=\frac{1}{2}\left[\rho\left(s|\Delta\right)\mathcal{L}+\mathcal{L}\rho\left(s|\Delta\right)\right].$ (29)
@@
$\mathcal{F}_{Q}\left(s|\Delta\right)\leq\mathcal{F}_{Q}\left(s\right),$ (30)
where the latter can be estimated via (27) through unitary evolution. Note that (30) can be readily obtained with the monotonicity of the Uhlmann-Jozsa fidelity of two quantum states $\rho$ and $\tilde{\rho}$ Nielsen and Chuang (2012), i.e.,
@@
Due to the geometric locality of the generator and of the quantum circuit applied to the system as well as the very small evolved time step, the generators $\{O_{i}\}$ that govern the time evolution of the subsystem $\Delta$ are restricted to include only those supported on $\Delta$ and its neighboring environment, where the latter is directly coupled to $\Delta$ by applying the quantum circuit. We denote $\tilde{\Delta}$ as the subsystem $\Delta$ and its neighboring environment, where the size $|\tilde{\Delta}|\leq\text{constant}$. Consequently, the summation on the right-hand side of (27) for upper bounding $\mathcal{F}_{Q}\left(s|\Delta\right)$ must be restricted to the basis operators which are locally supported on $\Delta$ and its neighboring environment, i.e.,
$\frac{1}{2}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\leq\frac{1}{2}\sqrt{\mathcal{F}_{Q}\left(s\right)}\leq\underset{i\ :\ \mathcal{S}_{i}\subset\tilde{\Delta}}{\sum}|h_{i}(s)|,$ (32)
@@
$\underset{\Delta}{\sum}\ \frac{1}{2}\int_{0}^{1}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\ \mathrm{d}s\leq\int_{0}^{1}\ \underset{\Delta}{\sum}\ \ \underset{i\ :\ \mathcal{S}_{i}\subset\tilde{\Delta}}{\sum}|h_{i}(s)|\ \mathrm{d}s\approx\int_{0}^{1}\underset{i}{\sum}|h_{i}(s)|\ \mathrm{d}s,$ (33)
@@
$\frac{1}{\sqrt{2}}D_{B}\left[\rho_{0}\left(\Delta\right),\rho_{1}\left(\Delta\right)\right]\leq\frac{1}{2}\int_{0}^{1}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\ \mathrm{d}s,$ (34)
which always holds for arbitrary smooth paths between the two reduced density matrices, caused by the smooth unitary path of the total system followed by the partial-trace operation, i.e., it is path-independent. Then we can take the infimum over all possible paths of the unitary QPP on the right-hand side of the last equation in (33) to obtain a tighter bound, and conclude that Nielsen’s QCC of geometrically local quantum circuits is approximately lower bounded by the summation over fidelity distances of the reduced density matrices supported on non-overlapping subsystems $\{\Delta\}$ which, together with their neighboring environments, cover the total system, i.e.,
$\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)\gtrsim\frac{1}{\sqrt{2}}\,\underset{\Delta}{\sum}\,D_{B}\left[\rho_{0}\left(\Delta\right),\rho_{1}\left(\Delta\right)\right].$ (35)
@@
Proof of Theorem  2: Here we follow a similar procedure as in Eisert (2021), but with more details in order to generalize the result to two arbitrary multi-qubit states in a generic QPP problem. Arrange the $n$-qubit system into a qubit chain $\{1,2,\cdots,n\}$ following a specific order, and take a subsystem over the $k$-th cut, $A_{k}=\{1,2,\cdots,k\}$, and its environment $B_{k}=\{k+1,k+2,\cdots,n\}$ as in the statement before Theorem 2. Further, we assume that the coupling in the generator $G(s)$ of $U(s)$ between $A_{k}$ and $B_{k}$ is spatially local with respect to the qubit chain. By Trotter decomposition, slicing $s\in[0,\,1]$ into small time steps $\{s_{j}\}_{j=0}^{T_{s}-1}$, and at each time interval $\Delta s_{j}=\left|s_{j}-s_{j-1}\right|$ ($j\in[1,T_{s}-1]$), the unitary evolution operator of the QPP is given by
$U(s_{j})=\mathrm{exp}\left[-i\Delta s_{j}G(s_{j})\right].$ (36)

### v2

@@ II.2 Unsupervised machine learning of topological order derived from quantum circuit complexity
For a family of smoothly parametrized quantum states $\rho(s)$ by $s\in[0,\,1]$, the quantum Fisher information (QFI) of the quantum state with respect to the parameter $s$ is given by Helstrom (1976); Holevo (1982); Ma et al. (2011); Taddei et al. (2013); Liu et al. (2020b) $\mathcal{F}_{Q}\left(s\right)=\mathrm{tr}\left[\rho\left(s\right)\mathcal{L}^{2}\right]$, with the symmetric logarithmic derivative (SLD) $\mathcal{L}$ determined by $\partial_{s}\rho\left(s\right)=\frac{1}{2}\left[\rho\left(s\right)\mathcal{L}+\mathcal{L}\rho\left(s\right)\right]$. The QFI, which plays a central role in quantum parameter estimation and which is also used for characterizing multipartite entanglement and topological states Zhang et al. (2018b), can be related to the infinitestmal Bures distance Braunstein and Caves (1994) with
@@
$\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\underset{h(s)}{\mathrm{inf}}\ \frac{1}{2}\int^{1}_{0}\sqrt{{\mathcal{F}}_{Q}(s)}\ \ \mathrm{d}s,$ (7)
which represents the optimal Bures distance between the two quantum states Taddei et al. (2013), reflecting the overall fidelity variation, as indicated by (6). Note that we have a symmetric QFC under permutations of its inputs, i.e., $\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\mathcal{C}_{\mathcal{F}}\left(\rho_{1}\rightarrow\rho_{0}\right)$.
@@
(Quantum circuit complexity upper bounds quantum Fisher complexity and Bures distance.) The QCC for the QPP from $\rho_{0}$ to $\rho_{1}$ is lower bounded by the QFC and the Bures distance,
@@
The proof of Theorem 1 is given in Sec. I of the Supplementary Information, and an asymptotic and tightness analysis of the bounds in Theorem 1 is provided in Sec. III. Note that, without the constraint of geometric locality of the quantum circuit, the overall Bures distance in (8) is equivalent to the $L_{2}$-norm distance between the two density matrices, which has been shown not to be a suitable distance measure for topological phases Che et al. (2020).
@@
[Quantum circuit complexity upper bounds entanglement-profile distance (A generalized version of Observation 1 in Eisert (2021)).] The QCC for the QPP from $\rho_{0}$ to $\rho_{1}$ with a geometrically local quantum circuit (with respect to the $n$-qubit chain) is lower bounded by
@@
The proof of Theorem 2 is given in Sec. I of the Supplementary Information, and an asymptotic and tightness analysis of the bounds in Theorem 2 is provided in Sec. III. Theorem 2 can be readily generalized to situations where the qubits are arranged in higher spatial dimensions and other geometries, provided that the quantum circuit is applied in a geometrically local manner accordingly. Nielsen’s QCC is then lower bounded by the entanglement change distributed over suitable cuts covering non-trivial bonds of the system.
@@
II.3 Relation to classical shadows
@@ II.4 Numerical experiments
II.4.1 The bond-alternating XXZ spin chain
@@
In Fig. 2, we present a comparison between the fidelity- and entanglement-based kernels for this model, respectively, together with the two-dimensional representation in the diffusion space from manifold learning of the diffusion map algorithm Coifman et al. (2005); Nadler et al. (2006); Rodriguez-Nieva and Scheurer (2019); Che et al. (2020); Scheurer and Slager (2020). The ground-state properties (fidelities and entanglements) are obtained from DMRG calculations of the model. Fixing $h_{0}=0$, $\delta=3$ and varying the value of $J_{2}/J_{1}$, we find that the unsupervised learning produces distinct clusters that can be associated with three phases (trivial, symmetry broken, and topological), in good agreement with independent calculations of topological invariants Pollmann and Turner (2012); Elben et al. (2020); Huang et al. (2022). More details can be found in the caption of Fig. 2. While we are ultimately aiming at the unsupervised learning of topological order in the presence of non-local quantum entanglement, Fig. 2 shows that the fidelity- and entanglement-based kernels also perform well when clustering symmetry-broken phases and symmetry-protected topological orders Elben et al. (2020) that are characterized by short-range entanglement.
@@
II.4.2 Kitaev’s toric code
@@
where $v$ denotes a vertex and $p$ is the face of the lattice, respectively; $A_{v}=\underset{j\in\mathrm{star}(v)}{\prod}X_{j}$ and $B_{p}=\underset{j\in\mathrm{boundary}(p)}{\prod}Z_{j}$ are the Kitaev’s star and face operators, respectively. The number of qubits for this model is $n=2L_{x}L_{y}$. The toric code ground state is gapped and exhibits area-law entanglement Kitaev and Preskill (2006). Therefore, for the purpose of a proof-of-principle demonstration and in order to obtain the entanglement profile for the kernel with a small system size, we use a MPS representation and DMRG to find a reasonably accurate solution for the ground state. For instance, numerical experiments show that, for a $4\times 4$ lattice with periodic boundary conditions in both dimensions, a maximal MPS bond dimension of $64$ is sufficient to produce accurate results.
In Fig. 3 we plot a comparison between the unsupervised machine learning of the toric code ground state (blue dots) and the random product states (RPS) (red dots) of qubits (i.e., random bit strings of spin ups and downs), with fidelity- and entanglement-based kernels, respectively. Both kernels perform well in clustering the topologically ordered quantum state and the trivial phase, while the entanglement-based one is more robust under the application of two-qubit Haar random unitaries to both the toric code and the RPS. This follows from the comparison between Fig. 3 (b) and (d), where the silhouette coefficient Skl (b), which takes the value one for best clustering, is used for illustrating the performance degradation of the clustering results in Fig. 3 (a) and (c), respectively, under Haar noises. The entanglement created by applying the two-qubit random circuit grows exponentially with the circuit depth. More details about the plot can be found in the caption of Fig. 3. For more discussions on the effects of hardware noises and applied random quantum circuits, we refer the reader to Sec. V of the Supplementary Information. In addition, in Sec. VI of the Supplementary Information (see Fig. S1), we also use metric-MDS (which does not require a kernel formulation) based on the Bures and entanglement-profile distance metrics in (9) and (12), respectively, for unsupervised manifold clustering of the toric code and the product state. Similar conclusions can be obtained from Fig. S1.
@@
II.4.3 The extended toric code
@@
III Discussion
In summary, we used quantum circuit complexity, a vital concept in quantum computation, to understand and to build viable distance measures and kernels for the unsupervised machine learning of topological order, targeting interpretability and generalizability. By resorting to Nielsen’s quantum circuit complexity, two easier-to-implement distance metrics based on fidelity and entanglement were proposed and their effectiveness numerically verified. Equipped with our topologically informational distance metrics, kernel-based and non-kernel manifold learning can be applied for interpretable phase clustering of quantum many-body systems.
The shadow kernel learning of Huang et al. (2022) can be naturally explained by our method derived from the Nielsen’s QCC, when the fidelity-based distance and classical shadow representation of the quantum state are used. The alternative approach based on the entanglement-entropy profile encodes more information about entanglement patterns in topological orders, and thus exhibits superior robustness against two-qubit random noises; moreover, they are more interpretable in terms of the relationship between topological quantum orders and long-range entanglement. Recent theoretical studies Vermersch et al. (2024a, b) show the promise of efficient measurement of many-body entropies, which indicates that the entanglement-based distance can be calculated efficiently from experimental data. This implies that our method can be readily applied to experimentally prepared multi-qubit quantum states; for instance, data obtained from near-term quantum computing hardware. When limited to SPT order and compared to the recent results in Wu et al. (2024) about neural-network learning topological orders of the bound-alternating XXZ model, our method is fully unsupervised and results in sharper phase boundaries in clustering quantum phases of this model while providing good interpretability.
@@ III Discussion
IV Methods
The pivot of our method is an informational distance metric for topological dissimilarities, leveraging Nielsen’s geometric QCC of mapping two quantum many-body states as a primitive distance metric for manifold learning. This can result in both good interpretability and generalizability for unsupervised machine learning of topological quantum order. The theoretical toolbox includes differential manifolds, quantum Fisher information, entanglement-growth rate, concept of topological equivalence, and quantum circuit computation, etc. By bridging conceptual power to practical implementation, Theorems 1 and 2 lead to approximate substitutes of the QCC, i.e., the Bures distance (fidelity change) and the entanglement-entropy distance, respectively, which preserve partial but important information of topological dissimilarity and which are intuitively more comprehensive and more accessible in simulations and measurements. As a comparison, we find that the entanglement-based manifold learning contains more topological information and performs better than the Bures-distance based one under Haar random noises. Regarding specific manifold-learning algorithms, we use the well-established framework with diffusion map Coifman et al. (2005); Nadler et al. (2006), kernel PCA Schölkopf et al. (1998); Mika et al. (1999), and metric-MDS Demaine et al. (2021); Skl (a) for demonstrations in the current paper. We believe that other manifold-learning algorithms such as the t-SNE Maaten and Hinton (2008) and the Isomap Tenenbaum et al. (2000), equipped with our topologically informational distances, will give similar results.
In the Supplementary Information, we provide, respectively, the details for the proofs of the theorems, a scaling and tightness analysis of the proved bounds, a complexity and scaling analysis of the method with respect to the system size (or the number of qubits $n$) and the sample size $N$, effects of noises and random circuits, and a (non-kernel) metric-MDS manifold learning of the toric-code.
V Acknowledgments
@@
SUPPLEMENTARY INFORMATION for
“Quantum circuit complexity and unsupervised machine learning of topological order”
Yanming Che,1, 2 Clemens Gneiting,2 Xiaoguang Wang,3 and Franco Nori1, 2
1Department of Physics, University of Michigan, Ann Arbor, Michigan 48109-1040, USA
2Center for Quantum Computing, RIKEN, Wako-shi, Saitama 351-0198, Japan
3Department of Physics, Zhejiang Sci-Tech University, Hangzhou 310018, China
January 15, 2026
I. Proof of Theorems ........................................................................................................................................................................I
II. Topological equivalence of gapped ground states and quantum circuit complexity of adiabatic quantum computation ........................................................................................................................................................................II
III. Asymptotic and tightness analysis of the bounds in Theorems 1 and 2 ........................................................................................................................................................................III
IV. Complexity and scaling analysis of the method ........................................................................................................................................................................IV
V. Effects of noise and random quantum circuits ........................................................................................................................................................................V
VI. Manifold learning of the toric code based on the metric-multidimensional scaling (metric-MDS) ........................................................................................................................................................................VI
I Proof of Theorems
@@
${\cal{F}}_{Q}(s)$ $=$ $4\left(\langle\partial_{s}\psi(s)|\partial_{s}\psi(s)\rangle-\left|\langle\partial_{s}\psi(s)|\psi(s)\rangle\right|^{2}\right)$ (S1)
@@
$\text{Var}\left[G(s)\right]\leq\frac{1}{4}\left[\lambda_{\text{max}}(s)-\lambda_{\text{min}}(s)\right]^{2}\leq\|G(s)\|^{2}_{\infty},$ (S2)
where $\lambda(s)$ is the eigenvalue of $G(s)$; and to obtain the last inequality, we have used $|a-b|\leq|a|+|b|$ for real numbers $a$ and $b$, and that, by definition, the operator norm of the generator $\|G(s)\|_{\infty}=|\lambda(s)|_{\text{max}}$. The first inequality in (S2) is saturated when $|\psi(s)\rangle=\frac{1}{\sqrt{2}}\left(|\lambda_{\text{max}}(s)\rangle+|\lambda_{\text{min}}(s)\rangle\right)$, i.e., the quantum state is an equal weight superposition of eigenstates with maximal and minimum eigenvalues of $G(s)$, respectively; and the last inequality is saturated when $\lambda_{\text{max}}(s)=-\lambda_{\text{min}}(s)>0$.
@@
$\|G(s)\|_{\infty}\leq\sum_{\sigma}|h_{\sigma}(s)|,$ (S3)
@@
$\sqrt{\mathcal{F}_{Q}(s)}\leq 2\sum_{\sigma}|h_{\sigma}(s)|$ (S4)
for all possible paths $h_{\sigma}(s)$ ($s\in[0,1]$). Then by integrating over $s\in[0,1]$ on both sides of (S4) followed by taking the infimum over all possible paths, the first inequality in (8) of the main text holds naturally, i.e.,
$\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)\leq{\cal{C}}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right).$ (S5)
The second inequality in (8) of the main text,
$\frac{D_{B}\left(\rho_{0},\rho_{1}\right)}{\sqrt{2}}$ $\leq$ ${\cal{C}}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right),$ (S6)
@@
$G(s)=\sum_{i}h_{i}(s)\ O_{i},$ (S7)
where $O_{i}$ is an orthonormal operator basis with operator norm $\|O_{i}\|_{\infty}\leq 1$, then, by replacing the summation in the right-hand side of (S4) to go over all orthonormal operator bases of the generator, (S4) also holds, i.e.,
$\sqrt{\mathcal{F}_{Q}(s)}\leq 2\sum_{i}|h_{i}(s)|.$ (S8)
Next, a suitable similarity measure for topological quantum order imposes the additional constraint of geometric locality of the QPP, which requires that $O_{i}$ is a geometrically local operator for all possible values of $i$, i.e., $O_{i}$ is supported on a constant number of neighboring particles. We consider a reduced density matrix $\rho\left(s|\Delta\right)$ supported on a subsystem $\Delta$ of constant size, undergoing a nonunitary evolution along the unitary QPP of the total system generated by $G(s)$. While the relation between the Bures distance and the QFI in (6) of the main text still holds for mixed quantum states with the Uhlmann-Jozsa fidelity, the expression of the QFI for $\rho\left(s|\Delta\right)$ with respect to $s$ is more complicated than (S1). Now the QFI has the general form given by the symmetric logarithmic derivative (SLD) $\mathcal{L}$, with
$\mathcal{F}_{Q}\left(s|\Delta\right)=\mathrm{tr}\left[\rho\left(s|\Delta\right)\mathcal{L}^{2}\right],$ (S9)
@@
$\partial_{s}\rho\left(s|\Delta\right)=\frac{1}{2}\left[\rho\left(s|\Delta\right)\mathcal{L}+\mathcal{L}\rho\left(s|\Delta\right)\right].$ (S10)
@@
$\mathcal{F}_{Q}\left(s|\Delta\right)\leq\mathcal{F}_{Q}\left(s\right),$ (S11)
where the latter can be estimated via (S8) through unitary evolution. Note that (S11) can be readily obtained with the monotonicity of the Uhlmann-Jozsa fidelity of two quantum states $\rho$ and $\tilde{\rho}$ Nielsen and Chuang (2012), i.e.,
@@
Due to the geometric locality of the generator and of the quantum circuit applied to the system as well as the very small evolved time step, the generators $\{O_{i}\}$ that govern the time evolution of the subsystem $\Delta$ are restricted to include only those supported on $\Delta$ and its neighboring environment, where the latter is directly coupled to $\Delta$ by applying the quantum circuit. We denote $\tilde{\Delta}$ as the subsystem $\Delta$ and its neighboring environment, where the size $|\tilde{\Delta}|\leq\text{constant}$. Consequently, the summation on the right-hand side of (S8) for upper bounding $\mathcal{F}_{Q}\left(s|\Delta\right)$ must be restricted to the basis operators which are locally supported on $\Delta$ and its neighboring environment, i.e.,
$\frac{1}{2}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\leq\frac{1}{2}\sqrt{\mathcal{F}_{Q}\left(s\right)}\leq\underset{i\ :\ \mathcal{S}_{i}\subset\tilde{\Delta}}{\sum}|h_{i}(s)|,$ (S13)
@@
$\underset{\Delta}{\sum}\ \frac{1}{2}\int_{0}^{1}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\ \mathrm{d}s\leq\int_{0}^{1}\ \underset{\Delta}{\sum}\ \ \underset{i\ :\ \mathcal{S}_{i}\subset\tilde{\Delta}}{\sum}|h_{i}(s)|\ \mathrm{d}s\approx\int_{0}^{1}\underset{i}{\sum}|h_{i}(s)|\ \mathrm{d}s,$ (S14)
@@
$\frac{1}{\sqrt{2}}D_{B}\left[\rho_{0}\left(\Delta\right),\rho_{1}\left(\Delta\right)\right]\leq\frac{1}{2}\int_{0}^{1}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\ \mathrm{d}s,$ (S15)
which always holds for arbitrary smooth paths between the two reduced density matrices, caused by the smooth unitary path of the total system followed by the partial-trace operation, i.e., it is path-independent. Then we can take the infimum over all possible paths of the unitary QPP on the right-hand side of the last equation in (S14) to obtain a tighter bound, and conclude that Nielsen’s QCC of geometrically local quantum circuits is approximately lower bounded by the summation over fidelity distances of the reduced density matrices supported on non-overlapping subsystems $\{\Delta\}$ which, together with their neighboring environments, cover the total system, i.e.,
$\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)\gtrsim\frac{1}{\sqrt{2}}\,\underset{\Delta}{\sum}\,D_{B}\left[\rho_{0}\left(\Delta\right),\rho_{1}\left(\Delta\right)\right].$ (S16)
@@
Proof of Theorem 2: Here we follow a similar procedure as in Eisert (2021), but with more details in order to generalize the result to two arbitrary multi-qubit states in a generic QPP problem. Arrange the $n$-qubit system into a qubit chain $\{1,2,\cdots,n\}$ following a specific order, and take a subsystem over the $k$-th cut, $A_{k}=\{1,2,\cdots,k\}$, and its environment $B_{k}=\{k+1,k+2,\cdots,n\}$ as in the statement before Theorem 2. Further, we assume that the coupling in the generator $G(s)$ of $U(s)$ between $A_{k}$ and $B_{k}$ is spatially local with respect to the qubit chain. By Trotter decomposition, slicing $s\in[0,\,1]$ into small time steps $\{s_{j}\}_{j=0}^{T_{s}-1}$, and at each time interval $\Delta s_{j}=\left|s_{j}-s_{j-1}\right|$ ($j\in[1,T_{s}-1]$), the unitary evolution operator of the QPP is given by
$U(s_{j})=\mathrm{exp}\left[-i\Delta s_{j}G(s_{j})\right].$ (S17)